OpenStudy (anonymous):

GIVING MEDAL find the sum of n terms of follwing 2/3 + 8/9 + 26/27 + 80/81+.... n terms

OpenStudy (anonymous):

each term is 1 - (1/3)^n so the sum could be expressed as sum as i goes from 1 to n of 1 - the sum as i goes from 1 to n of (1/3)^i which equals n - ∑[1 to n] (1/3)^i

OpenStudy (anonymous):

Can you explain further @Tonks

OpenStudy (anonymous):

the sum of powers on thirds is a(n) = (3^n-1)/2 In general, the sum of powers of 1/q is a(n) = (q^n-1)/(q-1)